Explain in quantitative terms the

circumstances under which the following

reaction can proceed.

What would the ratio of [product] to [reactant] have to be

for the reaction to proceed spontaneously?

Show your work.

Citrate

isocitrate



G'

° = +13.3 kJ/mol
Answer #1

Citrate --------> isocitrate

\DeltaG = -RTlnKeq

13300 = -8.314*298*2.303logKeq

logKeq   = -13300/5705.8483

              =-2.331

logKeq   = -2.331

   Keq    = 10-2.331   = 0.00466

   Keq    = [isocitate]/[citrate]

   0.00466 = [isocitate]/[citrate]